If sin2A = xsin2B prove that tan(A+B)/tan(A-B) = (x+1)/(x-1)

Grade:11

3 Answers

kumar
6 years ago
sin2A
kumar
6 years ago
sin2A/sin2B = x/1 and using componendo and dividendo we can write (sin2A + sin 2B)/(sin2A – sin2B) = (x+1)/(x-1) and then apply sinC + sinD and sin C – sin D formula .
alok yadav
4 years ago
bahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhodbahan ka land madharcvhod